# Deep Model Interpretation with Limited Data : A Coreset-based Approach

### v1

@@ 3.2 Coreset selection method modification
While our interpretation framework allows for any geometry-based coreset selection, we have considered two recent proposed coreset selection methods namely Moderate Selection [37] and D2Pruning [20].
D2Pruning. D2Pruning combines aspects of both geometry-based and un-certainty-based approaches. It constructs a graph of the dataset, where each node represents a data sample, and edges between nodes are determined by the Euclidean distance between their corresponding activation maps computed from the model. The initial score of each node is determined based on the collected training dynamics namely Area Under the Margin (AUM) which measures the probability gap between the target class and next largest class across all training epochs. The selection procedure is determined by updating the nodes’ score by forward and backward message transfer among nodes as formulated in [20].
@@ 3.3 Coreset-based evaluation protocol
Towards this goal, we compute activation maps $\boldsymbol{A}^{c}{\in}\mathbb{R}^{n^{c}\times w\times h\times d}$ from last convolutional layer, pertaining to all $n^{c}$ samples from class $c$. Given $\boldsymbol{A}^{c}$, we employ both $\boldsymbol{R}$ and $\boldsymbol{R^{\prime}}$ to compute interpretation maps $\boldsymbol{Z}^{c}$ and $\boldsymbol{Z}^{\prime c}$ with the shape $n^{c}{\times}w^{\prime}{\times}h^{\prime}{\times}d^{\prime}$. The computation of these interpretation maps is done according to the protocol of each interpretation method (Sec. 4.1). We reshape the interpretation maps into two dimensions $(n^{c}.w^{\prime}.h^{\prime}){\times}d^{\prime}$. Since the number of samples and data distributions between the coreset and the full dataset are different, there is no guarantee that interpretation elements in both $\boldsymbol{R}$ and $\boldsymbol{R^{\prime}}$ are aligned; the same holds between $\boldsymbol{Z}^{c}$ and $\boldsymbol{Z}^{\prime c}$. Hence, we consider an alignment-based similarity metric that finds an optimal transformation to align two representations.
@@ 5 Conclusion
Acknowledgements: This work is supported by the UAntwerp BOF DOCPRO4-NZ Project (id 41612) ”Multimodal Relational Interpretation for Deep Models”.

### v2

@@ 3.2 Coreset selection method modification
While our interpretation framework allows for any geometry-based coreset selection, we have considered two recent proposed coreset selection methods namely Moderate Selection [38] and our modification of D2Pruning [20].
DGPruning. D2Pruning combines aspects of both geometry-based and un-certainty-based approaches. It constructs a graph of the dataset, where each node represents a data sample, and edges between nodes are determined by the Euclidean distance between their corresponding activation maps computed from the model. The initial score of each node is determined based on the collected training dynamics namely Area Under the Margin (AUM) which measures the probability gap between the target class and next largest class across all training epochs. The selection procedure is determined by updating the nodes’ score by forward and backward message transfer among nodes as formulated in [20]. However, since our interpretation framework takes a pretrained model as input, we adapt this selection method to operate without the need of training dynamics. Therefore, drawing inspiration from [38], we determine the initial node scores based on the Euclidean distance of their activation maps to the class center. The class center is computed by averaging the activation maps belonging to the given class. The rest of the selection procedure remains unchanged. We refer to this modification as Diversity-Graph based Pruning (DGPruning).
@@ 3.3 Coreset-based evaluation protocol
Towards this goal, we compute activation maps $\boldsymbol{A}^{c}{\in}\mathbb{R}^{n^{c}\times w\times h\times d}$ from last convolutional layer, pertaining to all $n^{c}$ samples from class $c$. Given $\boldsymbol{A}^{c}$, we employ both $\boldsymbol{R}$ and $\boldsymbol{R^{\prime}}$ to compute interpretation maps $\boldsymbol{Z}^{c}$ and $\boldsymbol{Z}^{\prime c}$ with the shape $n^{c}{\times}w^{\prime}{\times}h^{\prime}{\times}d^{\prime}$. The computation of these interpretation maps is done according to the protocol of each interpretation method (Sec. 4.1). We reshape the interpretation maps into two dimensions $(n^{c}.w^{\prime}.h^{\prime}){\times}d^{\prime}$. Since the number of samples and data distributions between the coreset and the full dataset are different, there is no guarantee that interpretation elements in both $\boldsymbol{R}$ and $\boldsymbol{R^{\prime}}$ are aligned; the same holds between $\boldsymbol{Z}^{c}$ and $\boldsymbol{Z}^{\prime c}$. Hence, we consider an alignment-based similarity metric that finds an optimal transformation to align two representations. To do so, Partial-Whitening Shape Metric [36], denoted as $\phi(\bullet)$, is applied for measuring the angular distance between $\boldsymbol{Z}^{{}^{c}}$ and $\boldsymbol{Z^{\prime}}^{{}^{c}}$. This procedure is done for $C$ classes and similarity is reported as the average across classes (Eq.1). The angular distance is bounded in the interval $[0,\pi]$ [36]. As a result, a lower angular distance indicates highest similarity.
@@ 4.1 Interpretation framework setup
Interpretation maps computation. We employ the obtained relevant features, i.e., $\boldsymbol{R}$ and $\boldsymbol{R^{\prime}}$ by each interpretation method on the activation maps to compute interpretation maps $\boldsymbol{Z^{c}}$ and $\boldsymbol{Z^{\prime c}}$, respectively. This operation is performed according to the protocol of each interpretation method. Since ICE [40] and Topic-based interpretation [39] were originally applied to the last convolutional layer, the activation maps from this layer are modified by applying the learned relevant features. In the case of ICE, it reduces the depth of the activation maps using learned components (learned relevant features), followed by inverting the result into the activation maps dimensionality, considering as interpretation maps. In the case of Topic-based interpretation, activation maps are multiplied by the learned topics (learned relevant features). The resulted maps are considered as interpretation maps. Different from these, VEBI considers the activation maps from all layers and identifies indices of the convolutional filters, i.e. internal units, that are indicators of the classes of interest. Then, the concatenation of the activation maps per layer produced by the identified relevant units from the same layer are considered as the interpretation maps.